# LongStream: Long-Sequence Streaming Autoregressive Visual Geometry

### v1

@@ Abstract
Long-sequence streaming 3D reconstruction remains a significant open challenge. Existing autoregressive models often fail when processing long sequences. They typically anchor poses to the first frame, which leads to attention decay, scale drift, and extrapolation errors. We introduce LongStream, a novel gauge-decoupled streaming visual geometry model for metric-scale scene reconstruction across thousands of frames. Our approach is threefold. First, we discard the first-frame anchor and predict keyframe-relative poses. This reformulates long-range extrapolation into a constant-difficulty local task. Second, we introduce orthogonal scale learning. This method fully disentangles geometry from scale estimation to suppress drift. Finally, we solve Transformer cache issues such as attention-sink reliance and long-term KV-cache contamination. We propose cache-consistent training combined with periodic cache refresh. This approach suppresses attention degradation over ultra-long sequences and reduces the gap between training and inference. Experiments show LongStream achieves state-of-the-art performance. It delivers stable, metric-scale reconstruction over kilometer-scale sequences at 18 FPS. Project Page: https://3dagentworld.github.io/longstream/
Figure 1: Streaming Autoregressive Model Comparison for Metric-Scale Scene Reconstruction. Existing streaming models (e.g., Stream3R [20], StreamVGGT [69]) collapse within tens of meters, suffering from extrapolation errors. In contrast, our proposed LongStream delivers stable, kilometer-scale reconstruction. Its gauge-decoupled formulation and cache-consistent inference preserve metric accuracy and geometric stability, sustaining 18 FPS performance across multi-kilometer sequences.
@@ 1 Introduction
Geometry reconstruction, the joint estimation of camera poses and dense 3D structure from image sequences [36, 37, 4, 59, 40], is a cornerstone technology for applications like autonomous driving, AR/VR, and embodied robotics. These domains demand systems that can robustly process long-sequence video streams in real time.
Conventional pipelines [36, 37, 59, 60] and recent Transformer-based approaches [51, 21, 48, 52, 17] achieve state-of-the-art accuracy, but are inherently offline. They require reprocessing the entire sequence to integrate a new frame [20, 69], leading to massive computational redundancy and precluding real-time use. To address this, streaming models have been proposed. Recent works such as STream3R [20] and StreamVGGT [69] employ causal Transformers and KV-caching to build reconstructions incrementally. However, these models suffer from catastrophic extrapolation failure when processing long sequences. As shown in Figures 1 and  2, existing streaming methods incrementally reconstruct scenes in linear time, but their trajectories collapse within tens of meters, leading to complete tracking failure.
@@
  • Experiments across indoor and outdoor benchmarks show that LongStream attains state-of-the-art streaming reconstruction accuracy, while maintaining real-time throughput and stable metric scale on long sequences.
@@ 2 Related Work
Offline 3D reconstruction. Early end-to-end methods are mainly trained on image pairs [2, 55, 44, 64, 65]. Pointmaps offer efficient computation and real-time usage compared to voxel, mesh, or implicit fields [39, 15, 32, 28, 3], enabling SLAM and neural rendering [29, 19, 5, 6, 63]. DUSt3R [51] regresses pointmaps and relative pose from two images without intrinsics, but remains pairwise and requires global alignment. MASt3R [21] adds dense features and reciprocal matching to improve calibration and matching robustness, yet operates on pairs and needs costly fusion for multi-view scenes. VGGT [48] predicts poses, depth, pointmaps, and tracks from multi-view inputs in a single feed-forward pass, but relies on a fixed reference frame and absolute pose supervision, causing reference and scale biases. $\pi^{3}$ [52] removes reference-view bias via permutation-equivariant design and predicts local pointmaps, but outputs remain ambiguous up to a global similarity transform and lack metric scale.
@@ 3.3 Network Architecture
Following the design principles of VGGT [48], the architecture consists of a DINOv2-based [30] tokenizer, a causal Transformer aggregator, and task-specific heads for relative pose, depth, pointmap, and scale. Geometry is modeled in a keyframe-relative manner: each frame predicts its pose and pointmap with respect to the current keyframe $k$, enabling streaming reconstruction without reliance on a fixed first-frame coordinate frame.
@@
Scale head. To achieve $Sim(3)$ gauge invariance, we design an explicit scale head that receives the dedicated ScaleToken. It predicts an unconstrained log-scale variable $x_{s}\in\mathbb{R}$, which is then exponentiated to obtain a strictly positive scale factor:
@@ 3.4 Probabilistic Framework and Loss Functions
Table 1: Quantitative comparison on the KITTI dataset in terms of ATE. The upper block lists optimization-based baselines, and the lower block reports streaming methods. Our approach achieves the best accuracy across all sequences, with clear improvements on long-range trajectories.
@@ 3.5 KV Cache and Train–Inference Consistency
Prior work on streaming Transformers shows that models often rely on an attention sink [54, 56, 58] to stabilize attention; without it, sliding past the first frame can trigger model collapse.
@@ 4 Experiments
Figure 6: Qualitative comparison on indoor sequences. We evaluate challenging scenes with strong viewpoint changes, occlusions, and repeated back-tracking. While Stream3R, StreamVGGT, and VGGT-SLAM drift on these highly folded trajectories, LongStream maintains stable poses and consistent 3D structure throughout the sequence. * indicates OOM or repeated tracking loss.
@@ 4.1 Implementation Details
Training data. LongStream is trained on a multi-domain dataset collection, including Kubric [16], WildRGB [53], ScanNet [8], HyperSim [35], Mapillary [25], Replica [41], MVS-Synth [18], PointOdyssey [67], Virtual KITTI [1], Aria Synthetic Environments [31], Aria Digital Twin [31] Objaverse [11], Spring [27], and Waymo Open [43]. BlendedMVS [61], Co3Dv2 [34], MegaDepth [22], and DL3DV [23] do not provide metric ground truth, so we exclude them from scale training.
Baselines. We compare LongStream with offline transformers, streaming models, and SLAM-style systems. Both VGGT [48] and $\pi^{3}$ [52] cannot run on long clips due to memory limits, so we report FastVGGT [38] as a practical replacement. Streaming baselines include CUT3R [50], TTT3R, STream3R [20], and StreamVGGT [69], while MASt3R-SLAM [29] and VGGT-SLAM [26] serve as incremental SLAM counterparts. VGGT-SLAM performs windowed multi-frame inference per pass rather than frame-by-frame updates, so we treat it as an offline baseline. All baselines are run with the official default settings under a unified evaluation protocol.
@@ 4.2 Quantitative Results
Camera pose estimation. We evaluate ATE on vKITTI [1] (training), Waymo [43] (held-out), and the unseen KITTI [19], TUM-RGBD [42], and Oxford Spires [45] datasets. As shown in Tables 13, LongStream outperforms both offline and streaming baselines. While offline models suffer memory overflows (OOM) on long sequences, our method achieves state-of-the-art accuracy at $18$ FPS, demonstrating robust generalization across both large- and small-scale environments.
3D reconstruction. We evaluate full-sequence reconstruction on 7Scenes and TUM, reporting Chamfer Distance and F1@0.25. As shown in Table 4, LongStream performs competitively with both offline approaches on both benchmarks. The small spatial extent and extremely dense frame coverage of these datasets naturally compress numerical performance differences, making the metrics largely saturated. Nevertheless, as illustrated in Figure 6, our estimated camera trajectories exhibit noticeably higher stability over full sequences, revealing differences that saturated metrics fail to capture.
Scale estimation. We evaluate the accuracy of the recovered metric scale on vKITTI. LongStream produces a stable scale estimate across the entire sequence, achieving a scale ratio of $0.9905$ with respect to ground truth. Other streaming baselines do not provide accurate or temporally consistent metric-scale estimates.
@@ 4.3 Qualitative Results
Figures 5 and 6 visualize trajectories on kilometer-level and room-scale sequences, confirming stable pose prediction under both large and small spatial extents. In outdoor settings (Figure 5), existing streaming methods such as STream3R and StreamVGGT suffer from accumulated drift over long trajectories, while the optimization-based VGGT-SLAM encounters memory limitations (OOM) on longer sequences. In contrast, LongStream preserves trajectory continuity and metric accuracy across several hundred meters, successfully closing large loops without explicit loop closure modules.
Similarly, in indoor environments (Figure 6), the model demonstrates robustness against highly folded camera trajectories characterized by strong viewpoint changes, occlusions, and repeated back-tracking. Where baselines tend to produce unstable or drift-prone poses under these erratic motion patterns, LongStream maintains a coherent global trajectory and reconstructs consistent 3D structure throughout the entire sequence.
@@ 5 Conclusion
LongStream delivers stable, metric-scale reconstruction over ultra-long sequences, overcoming the drift and extrapolation failures of existing methods. Its gauge-decoupled pose design and cache-consistent training preserve consistent geometry and scale across thousands of frames.
Limitation. The model still assumes a largely static world, relies on a heuristic keyframe schedule, and shows mild degradation in point-map consistency over very long windows. These limitations suggest clear directions for improving robustness and generality in future work.

### v2

@@ Abstract
Long-sequence streaming 3D reconstruction remains a significant open challenge. Existing autoregressive models often fail when processing long sequences because they anchor poses to the first frame, leading to attention decay, scale drift, and extrapolation errors. We introduce LongStream, a novel gauge-decoupled streaming visual geometry model for metric-scale scene reconstruction across thousands of frames under a strictly online, future-invisible setting. Our approach is threefold. First, we discard the first-frame anchor and predict keyframe-relative poses. This reformulates long-range extrapolation into a constant-difficulty local task. Second, we introduce orthogonal scale learning. This method fully disentangles geometry from scale estimation to suppress drift. Finally, we identify attention bias issues in Transformers, including attention-sink reliance and long-term KV-cache saturation. We propose cache-consistent training combined with periodic cache refresh. This approach suppresses attention biases and contamination over ultra-long sequences and reduces the gap between training and inference. Experiments show that LongStream achieves state-of-the-art performance, enabling stable, metric-scale reconstruction over kilometer-scale sequences at 18 FPS. Project Page: https://3dagentworld.github.io/longstream/
Figure 1: Streaming Autoregressive Model Comparison for Metric-Scale Scene Reconstruction. Existing streaming models (e.g., Stream3R [21], StreamVGGT [75]) collapse within tens of meters, suffering from extrapolation errors. In contrast, our proposed LongStream delivers stable, kilometer-scale reconstruction. Its gauge-decoupled formulation and cache-consistent training preserve metric accuracy and geometric stability, sustaining 18 FPS performance across multi-kilometer sequences.
@@ 1 Introduction
Geometry reconstruction, the joint estimation of camera poses and dense 3D structure from image sequences [37, 38, 4, 65, 44, 54], is a cornerstone technology for applications like autonomous driving, AR/VR, and embodied robotics [39, 40]. These domains demand systems that can robustly process long-sequence video streams in real time.
Conventional pipelines [37, 38, 65, 66, 53] and recent Transformer-based approaches [57, 22, 51, 58, 17, 52] achieve state-of-the-art accuracy, but are inherently offline. They require reprocessing the entire sequence to integrate a new frame [21, 75], leading to massive computational redundancy and precluding real-time use. To address this, streaming models have been proposed. Recent works such as STream3R [21] and StreamVGGT [75] employ causal Transformers and KV-caching to build reconstructions incrementally. However, these models suffer from catastrophic extrapolation failure when processing long sequences. As shown in Figures 1 and  2, existing streaming methods incrementally reconstruct scenes in linear time, but their trajectories collapse within tens of meters, leading to complete tracking failure.
@@
  • Experiments across indoor and outdoor benchmarks show that LongStream achieves state-of-the-art performance under a strictly online, future-invisible setting, while maintaining real-time throughput and stable metric scale on long sequences.
@@ 2 Related Work
Offline 3D reconstruction. Early end-to-end methods are mainly trained on image pairs [2, 61, 70, 71]. Pointmaps offer efficient computation and real-time usage compared to voxel, mesh, or implicit fields [43, 15, 33, 29, 3], enabling SLAM and neural rendering [30, 20, 5, 6, 69]. DUSt3R [57] regresses pointmaps and relative pose from two images without intrinsics, but remains pairwise and requires global alignment. MASt3R [22] adds dense features and reciprocal matching to improve calibration and matching robustness, yet operates on pairs and needs costly fusion for multi-view scenes. VGGT [51] predicts poses, depth, pointmaps, and tracks from multi-view inputs in a single feed-forward pass, but relies on a fixed reference frame and absolute pose supervision, causing reference and scale biases. $\pi^{3}$ [58] removes reference-view bias via permutation-equivariant design and predicts local pointmaps, but outputs remain ambiguous up to a global similarity transform and lack metric scale.
@@ 3.3 Network Architecture
Following the architecture of VGGT-style streaming models [51, 21, 75], the architecture consists of a DINOv2-based [31] tokenizer, a causal Transformer aggregator, and task-specific heads for relative pose, depth, pointmap, and scale. Geometry is modeled in a keyframe-relative manner: each frame predicts its pose and pointmap with respect to the current keyframe $k$, enabling streaming reconstruction without reliance on a fixed first-frame coordinate frame.
@@
Scale head. To achieve $Sim(3)$ gauge invariance, we design an explicit scale head that receives a dedicated Scale Token, which is similar to the scale token idea in concurrent work MapAnything [19]. It predicts an unconstrained log-scale variable $x_{s}\in\mathbb{R}$, which is then exponentiated to obtain a strictly positive scale factor:
@@ 3.4 Probabilistic Framework and Loss Functions
Table 1: Quantitative comparison on the KITTI dataset in terms of ATE. The upper block lists optimization-based baselines, and the lower block reports streaming methods. Our approach achieves the best accuracy across all sequences.
@@ 3.5 KV Cache and Train–Inference Consistency
Prior work on streaming Transformers shows that models often rely on an attention sink [60, 62, 64, 42] to stabilize attention; without it, sliding past the first frame can trigger model collapse.
@@ 4 Experiments
Figure 6: Qualitative comparison on indoor sequences. We evaluate challenging scenes with strong viewpoint changes, occlusions, and repeated back-tracking. While Stream3R, StreamVGGT, and VGGT-SLAM drift on these highly folded trajectories, LongStream maintains stable poses and consistent 3D structure throughout the sequence. The symbol * denotes runs with out-of-memory errors or more than three tracking failures.
@@ 4.1 Implementation Details
Training data. LongStream is trained on a multi-domain dataset collection, including Kubric [16], WildRGB [59], ScanNet [8], HyperSim [36], Mapillary [26], Replica [45], MVS-Synth [18], PointOdyssey [73], Virtual KITTI [1], Aria Synthetic Environments [32], Aria Digital Twin [32] Objaverse [11], Spring [28], and Waymo Open [47]. BlendedMVS [67], Co3Dv2 [35], MegaDepth [23], and DL3DV [24] do not provide metric ground truth and are therefore excluded from scale training, while all other datasets used in training provide metric supervision.
Baselines. We compare LongStream with offline transformers, streaming models, and SLAM-style systems. Because both VGGT [51] and $\pi^{3}$ [58]exceed memory limits on long clips, we report FastVGGT [41] as a practical replacement. Streaming baselines include CUT3R [56], TTT3R, STream3R [21], and StreamVGGT [75]. MASt3R-SLAM [30] and VGGT-SLAM [27] represent optimization-based SLAM baselines. VGGT-SLAM performs chunk-wise inference over 16/32 frames and thus accesses future frames within each chunk. All methods are evaluated with official default settings under a unified protocol.
@@ 4.2 Quantitative Results
Camera pose estimation. We evaluate ATE on vKITTI [1] (training), Waymo [47] (held-out), and the unseen KITTI [20], TUM-RGBD [46], and Oxford Spires [48] datasets. As shown in Tables 13, streaming baselines exhibit nonlinear error growth as sequence length increases, likely due to long-horizon history saturation and contamination, whereas LongStream remains stable. Offline models frequently encounter out-of-memory (OOM) and tracking loss on long clips. Despite operating fully streaming, LongStream achieves state-of-the-art accuracy at $18$ FPS and generalizes robustly across environments.
3D reconstruction. We evaluate full-sequence reconstruction on 7Scenes and TUM on test sequences without subsampling, reporting Chamfer Distance and F1@0.25. As shown in Table 4 and Figure 6, LongStream performs competitively with both offline approaches on both benchmarks. The small spatial extent and dense frame coverage of these datasets lead to largely saturated metrics. For 7Scenes, failure cases of some methods under all-frame evaluation produce CD values several orders of magnitude larger on certain scenes, leading to an inflated mean CD.
Scale estimation. We evaluate the accuracy of the recovered metric scale on vKITTI. LongStream produces a stable scale estimate across the entire sequence, achieving a scale ratio of $0.9905$ with respect to ground truth.
@@ 4.3 Qualitative Results
Figures 5 and 6 visualize trajectories on kilometer-level and room-scale sequences, confirming stable pose prediction under both large and small spatial extents. In outdoor settings (Figure 5), existing streaming methods such as STream3R and StreamVGGT suffer from accumulated drift over long trajectories, while the optimization-based VGGT-SLAM encounters memory limitations and tracking loss on longer sequences. In contrast, LongStream preserves trajectory continuity and metric accuracy across several hundred meters, successfully closing large loops without explicit loop closure modules.
Similarly, in indoor environments (Figure 6), the model demonstrates robustness against highly folded camera trajectories characterized by strong viewpoint changes, occlusions, and repeated back-tracking. Where baselines tend to produce unstable or drift-prone poses under these erratic motion patterns, LongStream maintains a coherent global trajectory.
@@ 5 Conclusion
LongStream delivers stable, metric-scale reconstruction over ultra-long sequences, overcoming the drift and extrapolation failures of existing methods. Its gauge-decoupled pose design and cache-consistent training preserve consistent geometry and scale across thousands of frames, achieving strong accuracy and real-time performance on diverse indoor and outdoor datasets.
Limitations. The model still assumes a largely static world, relies on a heuristic keyframe schedule, and shows mild degradation in pointmap consistency over very long windows. These limitations suggest clear directions for improving robustness and generality in future work.
6 Acknowledgments
This work is supported by the National Natural Science Foundation of China (No. 62406267), Guangdong Provincial Project (No. 2024QN11X072), Guangzhou-HKUST(GZ) Joint Funding Program (No. 2025A03J3956) and Guangzhou Municipal Education Project (No. 2024312122).
@@ 8 Additional Attention Visualization Analysis
9 Long-Sequence Stability Analysis
Table 6 reports long-sequence results on Waymo #520018670 (135 m) and KITTI #03 (561 m). Streaming methods, including baselines and LongStream without refresh, exhibit non-linear error growth as sequence length increases. This suggests that under strictly-online constraints, longer histories or revisiting do not necessarily improve performance and may instead amplify long-horizon effects as the sequence grows. In contrast, LongStream remains stable over long sequences by removing first-frame anchoring and mitigating long-history effects through cache-consistent training with periodic cache refresh.
Method 15x 30x 60x 199x 30x 801x
CUT3R 0.159 1.421 2.505 8.591 3.867 148.06
TTT3R 0.153 0.873 1.127 5.505 1.957 105.28
Stream3R 0.181 1.329 2.998 21.440 3.412 158.25
VGGT-SLAM 0.172 0.518 0.992 3.740 0.929 169.83
Ours (w/o SW) 0.151 0.192 0.477 1.699 0.347 20.83
Ours 0.151 0.183 0.343 0.723 0.164 3.81
Table 6: ATE (m) as sequence length increases on Waymo #520018670 (Left, 135 m) and KITTI #03 (Right, 561 m). w/o SW denotes the variant without cache refresh and sliding window.